# SAGI: Semantically Aligned and Uncertainty Guided AI Image Inpainting

### v2

@@ 4.3 Localization and Detection Results
SAGI-D CAT-Net 69.2 +29.9 23.0 +2.1 46.5 +5.1 36.2 +30.3 98.8 +39.0 53.0 +10.6 99.9 +15.4 99.9 +60.0
PSCC 58.1 +23.0 57.6 +33.6 43.2 +23.1 19.6 +10.4 63.4 +3.5 71.8 +18.8 50.1 +9.8 35.6 +2.2
MMFusion 63.7 +44.1 41.4 +22.3 73.1 +26.8 50.3 +33.4 92.8 +31.1 80.9 +17.5 85.9 +23.0 81.5 +56.0
TruFor 64.7 +52.5 65.5 +46.4 86.8 +46.8 72.4 +52.8 96.2 +38.0 92.8 +32.7 95.5 +54.5 94.4 +81.9
Table 4: Evaluation of the SAGI-D generative pipeline for training forgery detection and localization models. Columns indicate training data, with Mean IoU for localization and Accuracy for detection. Results include in-domain (ID) and out-of-domain (OOD) performance for spliced (SP) and fully-regenerated (FR) images. Green numbers show improvements over original models. Retraining on SAGI-D significantly enhances performance, underscoring the value of semantic alignment for high-quality synthetic data.
@@
Table 5: Performance comparison of image forensics methods CAT-Net (CN), MMFusion (MM), PSCC (PS), and TruFor (TF) across TGIF and SAGI-D datasets. Metrics include detection and localization AUC. Data indicates training source. Green and red numbers show performance improvements and decreases over original models. Retraining on SAGI-D yields more consistent improvements, even across datasets, compared to training in TGIF.
We evaluate SAGI-D by comparing the performance of four state-of-the-art inpainting detection models: PSCC, CAT-Net, TruFor, and MMFusion. We assess both pre-trained models and versions retrained on our dataset. Table 4 presents both the localization (IoU) and detection (accuracy) results for in-domain and out-of-domain testing sets and spliced (SP) and fully regenerated (FR) images.
Retraining on SAGI-D leads to significant performance improvements across all models. TruFor shows the largest gains, with IoU increasing by +52.5% (in-domain) and +46.4% (out-of-domain), while CAT-Net achieves the highest in-domain IoU (69.2%, +29.9% improvement). However, CAT-Net struggles with out-of-domain generalization, improving only +2.1%, highlighting a disconnect between high accuracy and precise localization. Fully regenerated (FR) regions remain challenging, especially for original models, which achieve IoUs between 5.9% and 19.6%. Retrained models show substantial improvements on FR tasks, with TruFor reaching 72.4% IoU (+52.8%). SP detection is comparatively easier, with retrained TruFor achieving 86.8% IoU (+46.8%). Notably, TruFor and PSCC maintain consistent performance across in-domain and out-of-domain evaluations, while CAT-Net and MMFusion exhibit variability, likely due to their initial designs (e.g., JPEG-specific artifacts and added input complexities). Overall, SAGI-D proves highly effective for improving forgery localization, with TruFor emerging as the most robust model across domains and an important benchmark given that the original models underperform.
To further assess the need for SAGI-D as an inpainting benchmark, we compare models trained on their original data, TGIF, and SAGI-D, evaluating them on both TGIF and SAGI-D test sets (modified to avoid data leakage). Tab. 5 shows localization and detection AUC results. Training on SAGI-D consistently improves performance on SAGI-D, outperforming TGIF-trained models. Notably, SAGI-D-trained models sometimes surpass TGIF-trained ones even on TGIF (78.3 detection AUC compared to 74.3 on average). For instance, CAT-Net trained on SAGI-D achieves +28.6 and +21.8 improvements in detection AUC on SAGI-D and TGIF, respectively, while TGIF-trained CAT-Net drops -5.6 on TGIF. Similarly, PSCC trained on SAGI-D achieves +44.6 and +20.2 improvements in localization AUC on SAGI-D and TGIF, outperforming TGIF-trained PSCC (+7 on TGIF). SAGI-D trained TruFor even achieves AUCs above 90 on both datasets. Original models perform better on TGIF than SAGI-D, highlighting SAGI-D’s greater complexity and utility as a benchmark. These results demonstrate SAGI-D’s diversity and superiority for training forgery detection models over TGIF.
@@ 5 Conclusions
In this work we proposed a model-agnostic framework for generating and evaluating high-quality inpaintings, based upon the key ideas of semantic alignment and uncertainty guidance. Looking ahead, our framework can benefit from advances in foundation models to further refine prompt generation and realism assessment. Future efforts should focus on using more advanced VLMs for improved realism estimation, and integrating human-in-the-loop evaluation for finer perceptual alignment. We hope our framework will drive progress in both enhancing inpainting quality and strengthening forensic tools, ensuring safer and more realistic applications of generative AI as its adoption grows.
@@
While any advances in generative AI enable the automation of tasks that previously required significant human effort, providing significant societal benefits, they also enable their potential malicious exploitation, such as in spreading misinformation, or manipulating visual evidence. Yet, the fact that our proposed framework can be used for significantly enhancing the robustness of image forensic tools, allows mitigating such misuse risks.
@@ Acknowledgments:
This work was supported by the Horizon Europe vera.ai project (grant no.101070093) and by the High Performance Computing infrastructure of the Aristotle University of Thessaloniki.
@@ 6 Implementation details
This section provides additional implementation details of our approach to ensure reproducibility.
6.1 Source of Authentic Data
Since RAISE [18] contains RAW images, we processed these images before using them for inpainting experiments. We utilized the RAISE dataset as established in [34].
@@ 6.2 Dataset Splits
As shown in Table 6, we structure our dataset to evaluate both in-domain performance and generalization to new data. For in-domain evaluation, we use COCO (60,000 randomly selected training images and nearly all 5,000 validation images for validation and testing) and RAISE (7,735 images processed with $\Phi_{\text{seg}}$, yielding 25,674 image-mask-model combinations through 1-7 masks or prompts per image, with derived images kept in the same split, as each image was inpainted up to 4 times only in this dataset). To test generalization, we create an out-of-domain testing split using OpenImages—a dataset not used during training—comprising 6,000 randomly selected test images. This split uses a different language model $\Theta_{\text{llm}}$ (Claude) than COCO and RAISE (ChatGPT), providing a way to evaluate how well models perform on both new data and different prompting approaches. Throughout our experiments, we refer to the COCO and RAISE test splits as in-domain and the OpenImages test split as out-of-domain.
@@ 6.3 SAOR configuration
The specific API endpoints used in our implementation were gpt-3.5-turbo [50] (as of June 2024) and claude-3-5-sonnet-20240620 [3]. The system prompt used for the LLMs in SAOR was configured as shown in Fig. 4. For double inpainting cases, where two objects needed to be sequentially modified, we used an adapted system prompt to select a second object and generate a prompt as shown if Fig. 4. For images designated for object removal, we used a simplified system prompt focused solely on object selection that is shown if Fig. 4.
@@
All LLM interactions were configured with hyperparameters including a temperature of 1.2 to encourage creative variations in the generated prompts, a top-p (nucleus sampling threshold) of 0.8, and a maximum token limit of 40 for prompt length. Some prompts from initial experiments, conducted without the maximum token restriction, were retained in our final dataset. The prefix ”Inpaint the masked area with…” was included in the system prompts to maintain a consistent format in the LLMs’ responses but was omitted from the actual saved prompts to avoid potential misinterpretation by diffusion models.
@@ 6.4 Inpainting Pipelines Configuration
The text-guided inpainting models support Stable Diffusion [59] by default, along with certain community versions. Specifically, HD-Painter [44] supports Stable Diffusion v1-5111https://huggingface.co/stable-diffusion-v1-5/stable-diffusion-v1-5, Stable Diffusion v2222https://huggingface.co/stable-diffusion-v1-5/stable-diffusion-v1-5, and DreamShaper v8333https://civitai.com/models/4384/dreamshaper. BrushNet [26] supports Stable Diffusion v1-5, Stable Diffusion XL [^4] https://huggingface.co/stabilityai/stable-diffusion-xl-base-1.0, DreamShaper v8, Realistic Vision [^5] https://civitai.com/models/4201/realistic-vision-v60-b1, epiCRealism [^6] https://civitai.com/models/25694/epicrealism, and JuggernautXL [^7] https://civitai.com/models/133005/juggernaut-xl. PowerPaint [93] combines Realistic Vision and BrushNet, while ControlNetInpaint [86] supports Stable Diffusion 1.5. Inpaint-Anything [82] supports Stable Diffusion v2.
@@ 6.5 UGDA Configuration
Verdict: (final verdict, answer ”Yes, it is realistic” or ”No, it is not realistic”)
@@
Verdict: (final verdict, answer ”First is more realistic” or ”Second is more realistic” or ”Both look realistic”)
@@ 6.6 Detailed Classification Process of UGDA
In all other response combinations, $\hat{I}_{i}$ is classified as undeceiving. This classification scheme captures cases where the VLM either consistently prefers the inpainted image or shows uncertainty in its assessment, all of which indicate potential deceptiveness in the synthetic content. Case 1 indicates model uncertainty manifested through order sensitivity, Case 2 represents clear preference for synthetic content, Case 3 captures uncertainty biased toward synthetic content, and Case 4 reflects complete inability to distinguish between real and synthetic content. These patterns suggest varying degrees of image deceptiveness that warrant classification as deceiving.
@@ 6.7 Human Benchmark
The application implementing the human benchmark was developed using Gradio [1]. At the start of the demo, participants were provided with clear instructions on how to proceed. They were asked to evaluate whether an image had been inpainted and to draw bounding boxes around the areas they believed to be inpainted. Additionally, participants were asked to complete a short demographics questionnaire before beginning the task. The questions included are shown in Tab. 7:
@@ 8 Localization and Detection results
Since mean IoU and detection Accuracy require a threshold, we also report AUC metrics in Table 10 as they are threshold-agnostic. We calculate AUC at both pixel level (localization) and image level (detection). For localization AUC, we resize and flatten all localization maps and their ground truths into two vectors for ROC computation in each group. Note that detection AUC cannot be calculated for SP and FR sets, as they contain only forged images. The AUC metrics further confirm that retraining improves performance significantly. TruFor’s localization AUC increases from 68.9% to 99.5% for in-domain and 79.9% to 99.6% for out-of-domain testing. Similarly, CAT-Net shows strong in-domain gains (60.0% to 95.6%) but smaller out-of-domain improvement (51.7% to 90.8%). Domain generalization varies across models. While retrained CAT-Net achieves high in-domain detection AUC (99.6%), it drops to 76.for out-of-domain. In contrast, retrained TruFor maintains consistent performance across domains in both localization (99.5%/99.6%) and detection (99.2%/98.0%). SP localization remains easier for both original and retrained models than FR, with all retrained models achieving localization AUCs above 90.0% for SP tasks. For FR images, original models perform poorly (AUCs 52.7%-74.1%) but show clear improvements after retraining, with TruFor reaching 98.6% AUC.
@@
Table 10: Performance comparison of image forensics methods across different domains. The table shows AUC scores for both localization and detection tasks, comparing original models with those retrained on our dataset. ”ID” indicates in-domain and ”OOD” indicates out-of-domain performance, while SP (Splicing) and FR (Fully Regenerated) represent different forgery types. Green numbers show improvements and red numbers show decreases compared to original models. Retraining on our dataset yields significant performance improvements across most metrics and models.
@@
Table 11: Performance comparison of image forensics methods CAT-Net (CN), MMFusion (MM), PSCC (PS), and TruFor (TF) for both FR (Fully Regenerated) and SP (Splicing) scenarios. For each method, Mean IoU and Accuracy (Acc) scores are shown. Inpainting Models: BN (BrushNet), CN (ControlNet), HDP (HDPainter), PPt (PowerPaint), IA (InpaintAnything), and RA (RemoveAnything). The green values in the second row for each retrained model indicate the improvement over the corresponding original model.
@@ 8.2 Model Compression Robustness Analysis
We evaluate model robustness against JPEG and WEBP image compression at quality levels $0.85$, $0.7$, and $0.5$. Figure 7 presents detection and localization results. Retrained TruFor shows the strongest resilience, maintaining stable performance across quality levels for both compression types. WEBP compression affects performance more than JPEG, particularly for localization tasks. All models show higher degradation in IoU scores compared to accuracy metrics, indicating that manipulation localization is more sensitive to compression artifacts than detection.
@@ 8.3 Comparison with human performance
Table 13 presents the comparison between human evaluators and automated models. Human performance reached 67.4% accuracy and 15.2 IoU, significantly lower than retrained models like TruFor (95.3% accuracy, 68.3%) and MMFusion (87.9% accuracy, 50.9% IoU). Results are broken down into four categories: All represents performance on the complete test set, while Deceiving, Undeceiving and Intermediate correspond to UGDA’s classification of images based on their potential to fool human perception. The Intermediate category includes images that passed the initial realism check but not the second. Users particularly struggled with deceiving images (35.2% accuracy, 12.9% IoU) compared to undeceiving ones (73.7% accuracy, 40.4% IoU), validating UGDA’s effectiveness in identifying manipulations that are challenging for human perception. Also, the performance of humans on the intermediate category (59.7% accuracy, 28.3% IoU) confirms that the second stage is indeed efficient in discarding images that are not truly deceiving. In contrast, retrained models maintain high performance even on these challenging cases, with TruFor achieving 98.9% accuracy and 87.4% IoU on deceiving images. The performance gap between humans and models emphasizes the importance of automated detection methods, particularly for high-quality inpainting that can bypass human perception.
@@
Table 13: Human vs. model performance comparison on inpainting detection. Results show accuracy and IoU for full test set (All) and images classified by UGDA as Deceiving (Dec.) or Undeceiving (Und.). † indicates models retrained on our dataset. Bold values indicate the best performance per column.
In Table 14 we see the performance comparison between spliced (SP) and fully regenerated (FR) images, across human evaluators and the forensic models. The results reveal that human performance remains consistent across both manipulation types, showing no significant advantage in detecting either SP (0.34 for Deceiving, 0.76 for Undeceiving) or FR manipulations (0.37 for Deceiving, 0.74 for Undeceiving) in contrast to the forensic models.
@@
Table 14: Human vs. model performance comparison on inpainting detection. Results show accuracy and IoU for full test set images classified by UGDA as Deceiving (Dec.) or Undeceiving (Und.), SP and FR. † indicates models retrained on our dataset. Bold values indicate the best performance per column.
8.4 Qualitative Analysis of Localization Performance
In Figure 8, we present a comparison of the localization maps before and after fine-tuning. The results demonstrate that fine-tuning can significantly improve localization performance. For PSCC-Net, while improvements are observed in the second and fifth rows, the poor localization results persist in other cases. Regarding the remaining models, localization improvements are evident in all cases, with TruFor consistently demonstrating the most accurate localization maps. The second row showcases an example where the original CAT-Net, MMFusion, and TruFor successfully identified the inpainted area, while the fourth row presents a case where the original model could only partially detect the inpainted region. The fifth row presents an interesting case involving the original MMFusion model. If the predicted mask were inverted, it would have successfully identified the inpainted area. This can be attributed to the fact that in splicing it can be ambiguous which area is spliced and which is original. In AI inpainting cases, however, there is no ambiguity about which region has been modified.

### v3

@@ 4.3 Localization and Detection Results
SAGI-D
CAT-Net[36] 69.2 +29.9 23.0 +2.1 46.5 +5.1 36.2 +30.3 98.8 +39.0 53.0 +10.6 99.9 +15.4 99.9 +60.0
PSCC-Net[41] 58.1 +23.0 57.6 +33.6 43.2 +23.1 19.6 +10.4 63.4 +3.5 71.8 +18.8 50.1 +9.8 35.6 +2.2
MMFusion[75] 63.7 +44.1 41.4 +22.3 73.1 +26.8 50.3 +33.4 92.8 +31.1 80.9 +17.5 85.9 +23.0 81.5 +56.0
TruFor[24] 64.7 +52.5 65.5 +46.4 86.8 +46.8 72.4 +52.8 96.2 +38.0 92.8 +32.7 95.5 +54.5 94.4 +81.9
Table 5: Evaluation of SAGI on training image forgery detection and localization models. The first column indicates training data, while Mean IoU is reported for localization and Accuracy for detection. Results include in-domain (ID) and out-of-domain (OOD) performance for spliced (SP) and fully-regenerated (FR) images. ID and OOD performance is computed across inpainted and authentic images, while SP and FR highlight the performance on the corresponding type of inpaintings. Green numbers show improvements over original models. Retraining on SAGI-D significantly enhances performance, underscoring the value of semantic alignment for high-quality synthetic data.
@@
Table 6: Performance comparison of image forensics methods CAT-Net (CN), MMFusion (MM), PSCC-Net (PS), and TruFor (TF) across TGIF and SAGI-D datasets. Metrics include detection and localization AUC. Data indicates training source. Green and red numbers show performance improvements and decreases over original models. Retraining on SAGI-D yields more consistent improvements, even across datasets, compared to training on TGIF.
Figure 4: Failure cases for SAOR and UGDA. Top row shows SAOR generating semantically inappropriate content (a colorful hot air balloon floating in the sky). Bottom row demonstrates UGDA’s failure: 1st assessment gets fixated on “unnatural duplication of couch” without recognizing the blur, while 2nd assessment correctly identifies “couch appears altered with blurry texture”.
We evaluate SAGI-D using four state-of-the-art inpainting detection models: PSCC-Net, CAT-Net, TruFor, and MMFusion. We assess both pre-trained models and versions retrained on our dataset. Table 5 presents both the localization (IoU) and detection (accuracy) results for in-domain and out-of-domain testing sets and SP and FR images. Retraining on SAGI-D leads to significant performance improvements across all models. TruFor shows the largest gains, with IoU increasing by +52.5% (in-domain) and +46.4% (out-of-domain), while CAT-Net achieves the highest in-domain IoU (69.2%, +29.9% improvement). FR regions remain challenging for original models (IoUs between 5.9% and 19.6%), but retrained models show substantial improvements, with TruFor reaching 72.4% IoU (+52.8%). SP detection is easier, with retrained TruFor achieving 86.8% IoU. TruFor and PSCC-Net maintain consistent cross-domain performance, while CAT-Net and MMFusion exhibit variability due to their designs (e.g., JPEG-specific artifacts and added input complexities). Overall, SAGI-D proves highly effective for improving image forgery detection and localization models.
To further assess SAGI-D as a benchmark, we compare models trained on their original data, TGIF and SAGI-D, and evaluate them on both TGIF and SAGI-D test sets. Tab. 6 shows localization and detection AUC results. Training on SAGI-D consistently improves performance, outperforming TGIF-trained models. Notably, SAGI-D-trained models sometimes surpass TGIF-trained ones even on TGIF (78.3 vs 74.3 detection AUC average). CAT-Net trained on SAGI-D achieves +28.6 and +21.8 improvements on SAGI-D and TGIF respectively, while TGIF-trained CAT-Net drops -5.6 on TGIF. PSCC-Net shows +44.6 and +20.2 localization AUC gains, and SAGI-D-trained TruFor reaches AUCs above 90 on both datasets. Original models perform better on TGIF than SAGI-D, highlighting its greater complexity and benchmark value. These results affirm the superiority of SAGI-D for training detection models.
@@ 5 Conclusions
In this work we proposed a model-agnostic framework for generating and evaluating high-quality inpaintings, based on the key ideas of semantic alignment and uncertainty guidance. Looking ahead, our framework can benefit from advances in foundation models to further refine prompt generation and realism assessment and extend uncertainty mechanisms to enhance LLM/VLM performance in tasks involving reference samples, such as image-to-image translation, audio/video editing, and quality assessment of generated content. We aim for this framework to advance both generative quality and forensic detection, supporting safer applications of generative AI.
@@
While advances in generative AI provide significant societal benefits by automating previously labor-intensive tasks, they also enable potential malicious exploitation, such as spreading misinformation, or manipulating visual evidence. Yet, the fact that our proposed framework can significantly enhance the robustness of image forensic tools, allows mitigating such misuse risks.
@@ Acknowledgments:
This work was supported by the Horizon Europe projects vera.ai (grant no. 101070093), ELLIOT (101214398) and AI4TRUST (101070190). Compute resources were granted by GRNET and the HPC infrastructure of the Aristotle University of Thessaloniki.
@@ 6 Implementation details
This section provides additional implementation details of our approach to ensure reproducibility. Code is available on https://github.com/mever-team/SAGI.
6.1 Source of Authentic Images
Since RAISE [22] contains RAW images, we processed these images before using them for inpainting experiments. We utilized the RAISE dataset as described in [36].
@@ 6.2 Dataset Splits
As shown in Table 7, we structure our dataset to evaluate both in-domain performance and generalization to new data. For in-domain evaluation, we use COCO (60,000 randomly selected training images and nearly all 5,000 validation images for validation and testing) and RAISE (7,735 images processed with $\Phi_{\text{seg}}$, yielding 25,674 image-mask-model combinations through 1-7 masks or prompts per image, with derived images kept in the same split, as each image was inpainted up to 4 times only in this dataset). To test generalization, we create an out-of-domain testing split using OpenImages [8]—a dataset not used during training—comprising 6,000 randomly selected test images. This split uses a different language model $\Theta_{\text{llm}}$ (Claude) than COCO and RAISE (ChatGPT), providing a way to evaluate how well models perform on both new data and different prompting approaches. Throughout our experiments, we refer to the COCO and RAISE test splits as in-domain and the OpenImages test split as out-of-domain.
@@ 6.3 SAOR configuration
The specific API endpoints used in our implementation were gpt-3.5-turbo [53] (as of June 2024) and claude-3-5-sonnet-20240620 [4]. The system prompt used for the LLMs in SAOR was configured as shown in Fig. 5. For double inpainting cases, where two objects needed to be sequentially modified, we used an adapted system prompt to select a second object and generate a prompt as shown in Fig. 5. For images designated for object removal, we used a simplified system prompt focused solely on object selection that is shown in Fig. 5.
@@
All LLM interactions were configured with hyperparameters including a temperature of 1.2 to encourage creative variations in the generated prompts, a top-p (nucleus sampling threshold) of 0.8, and a maximum token limit of 40 for prompt length. Some prompts from initial experiments, conducted without the maximum token restriction, were retained in our final dataset. The prefix “Inpaint the masked area with…” was included in the system prompts to maintain a consistent format in the LLMs’ responses but was omitted from the actual saved prompts to avoid potential misinterpretation by diffusion models.
@@ 6.4 Inpainting Pipelines Configuration
The text-guided inpainting models support Stable Diffusion [62] by default, along with certain community versions. Specifically, HD-Painter [46] supports Stable Diffusion v1.5111https://huggingface.co/stable-diffusion-v1-5/stable-diffusion-v1-5, Stable Diffusion v2222https://huggingface.co/stabilityai/stable-diffusion-2, and DreamShaper v8333https://civitai.com/models/4384/dreamshaper. BrushNet [30] supports Stable Diffusion v1.5, Stable Diffusion XL [^4] https://huggingface.co/stabilityai/stable-diffusion-xl-base-1.0, DreamShaper v8, Realistic Vision [^5] https://civitai.com/models/4201/realistic-vision-v60-b1, epiCRealism [^6] https://civitai.com/models/25694/epicrealism, and JuggernautXL [^7] https://civitai.com/models/133005/juggernaut-xl. PowerPaint [97] combines Realistic Vision and BrushNet, while ControlNetInpaint [90] supports Stable Diffusion v1.5. Inpaint-Anything [86] supports Stable Diffusion v2.
@@ 6.5 UGDA Configuration
Verdict: (final verdict, answer “Yes, it is realistic” or “No, it is not realistic”)
@@
Verdict: (final verdict, answer “First is more realistic” or “Second is more realistic” or “Both look realistic”)
@@ 6.6 Detailed Classification Process of UGDA
In all other response combinations, $\hat{I}_{i}$ is classified as non-deceiving. This classification scheme captures cases where the VLM either consistently prefers the inpainted image or shows uncertainty in its assessment, all of which indicate potential deceptiveness in the synthetic content. Case 1 indicates model uncertainty manifested through order sensitivity, Case 2 represents clear preference for synthetic content, Case 3 captures uncertainty biased toward synthetic content, and Case 4 reflects complete inability to distinguish between real and synthetic content. These patterns suggest varying degrees of image deceptiveness that warrant classification as deceiving.
@@ 6.7 Human Benchmark
The application implementing the human benchmark was developed using Gradio [1]. At the start of the demo, participants were provided with clear instructions on how to proceed. They were asked to evaluate whether an image had been inpainted and to draw bounding boxes around the areas they believed to be inpainted. Additionally, participants were asked to complete a short demographics questionnaire before beginning the task. The questions included are shown in Tab. 8.
@@ 8 Localization and Detection Results
Since mean IoU and detection Accuracy require a threshold, we also report AUC metrics in Table 11 as they are threshold-agnostic. We calculate AUC at both pixel level (localization) and image level (detection). For localization AUC, we resize and flatten all localization maps and their ground truths into two vectors for ROC computation in each group. Note that detection AUC cannot be calculated for SP and FR sets, as they contain only forged images. The AUC metrics further confirm that retraining improves performance significantly. TruFor’s localization AUC increases from 68.9% to 99.5% for in-domain and 79.9% to 99.6% for out-of-domain testing. Similarly, CAT-Net shows strong in-domain gains (60.0% to 95.6%) but smaller out-of-domain improvement (51.7% to 90.8%). Domain generalization varies across models. While retrained CAT-Net achieves high in-domain detection AUC (99.6%), it drops to 76% for out-of-domain. In contrast, retrained TruFor maintains consistent performance across domains in both localization (99.5%/99.6%) and detection (99.2%/98.0%). SP localization remains easier for both original and retrained models than FR, with all retrained models achieving localization AUCs above 90.0% for SP tasks. For FR images, original models perform poorly (AUCs 52.7%-74.1%) but show clear improvements after retraining, with TruFor reaching 98.6% AUC.
@@
Table 11: Performance comparison of image forensics methods across different domains. The table shows AUC scores for both localization and detection tasks, comparing original models with those retrained on our dataset. “ID” indicates in-domain and “OOD” indicates out-of-domain performance, while SP (Splicing) and FR (Fully Regenerated) represent different forgery types. Green numbers show improvements and red numbers show decreases compared to original models. Retraining on our dataset yields significant performance improvements across most metrics and models.
@@
Table 12: Performance comparison of image forensics methods CAT-Net (CN), MMFusion (MM), PSCC-Net (PS), and TruFor (TF) for both FR (Fully Regenerated) and SP (Splicing) scenarios. For each method, Mean IoU and Accuracy (Acc) scores are shown. Inpainting Models: BN (BrushNet), CN (ControlNet), HDP (HDPainter), PPt (PowerPaint), IA (InpaintAnything), and RA (RemoveAnything). The green values in the second row for each retrained model indicate the improvement over the corresponding original model.
@@ 8.2 Model Compression Robustness Analysis
We evaluate model robustness against JPEG and WEBP image compression at quality levels 0.850.850.85, 0.70.70.7, and 0.50.50.5. Figure 8 presents detection and localization results. Retrained TruFor shows the strongest resilience, maintaining stable performance across quality levels for both compression types. WEBP compression affects performance more than JPEG, particularly for localization tasks. All models show higher degradation in IoU scores compared to accuracy metrics, indicating that manipulation localization is more sensitive to compression artifacts compared with detection.
@@ 8.3 Comparison with human performance
Table 14 presents the comparison between human evaluators and automated models. Human performance reached 67.4% accuracy and 15.2 IoU, significantly lower than retrained models like TruFor (95.3% accuracy, 68.3%) and MMFusion (87.9% accuracy, 50.9% IoU). Results are broken down into four categories: All represents performance on the complete test set, while Deceiving, Non-Deceiving and Intermediate correspond to UGDA’s classification of images based on their potential to fool human perception. The Intermediate category includes images that passed the initial realism check but not the second. Users particularly struggled with deceiving images (35.2% accuracy, 12.9% IoU) compared to non-deceiving ones (73.7% accuracy, 40.4% IoU), validating UGDA’s effectiveness in identifying manipulations that are challenging for human perception. Also, the performance of humans on the intermediate category (59.7% accuracy, 28.3% IoU) confirms that the second stage is indeed effective in discarding images that are not truly deceiving. In contrast, retrained models maintain high performance even on these challenging cases, with TruFor achieving 98.9% accuracy and 87.4% IoU on deceiving images. The performance gap between humans and models emphasizes the importance of automated detection methods, particularly for high-quality inpainting that can bypass human perception.
@@
Table 14: Human vs. model performance comparison on inpainting detection. Results show accuracy and IoU for full test set (All) and images classified by UGDA as Deceiving (Dec.) or Non-Deceiving (ND.). † indicates models retrained on our dataset. Bold values indicate the best performance per column.
In Table 15 we see the performance comparison between spliced (SP) and fully regenerated (FR) images, across human evaluators and the forensic models. The results reveal that human performance remains consistent across both manipulation types, showing no significant advantage in detecting either SP (0.34 for Deceiving, 0.76 for Non-Deceiving) or FR manipulations (0.37 for Deceiving, 0.74 for Non-Deceiving) in contrast to the forensic models.
@@
Table 15: Human vs. model performance comparison on inpainting detection. Results show accuracy and IoU for full test set images classified by UGDA as Deceiving (Dec.) or Non-Deceiving (Non-Dec.), SP and FR. † indicates models retrained on our dataset. Bold values indicate the best performance per column.
8.4 Qualitative Analysis
In Figure 9, we present a comparison of the localization maps before and after fine-tuning. The results demonstrate that fine-tuning can significantly improve localization performance. For PSCC-Net, while improvements are observed in the second and fifth rows, poor localization results persist in other cases. Regarding the remaining models, localization improvements are evident in all cases, with TruFor consistently demonstrating the most accurate localization maps. The second row showcases an example where the original CAT-Net, MMFusion, and TruFor successfully identified the inpainted area, while the fourth row presents a case where the original model could only partially detect the inpainted region. The fifth row presents an interesting case involving the original MMFusion model. If the predicted mask were inverted, it would have successfully identified the inpainted area. This can be attributed to the fact that in splicing it can be ambiguous which area is spliced and which is original. In AI inpainting cases, however, there is no ambiguity about which region has been modified.